# Know You First and Be You Better: Modeling Human-Like User Simulators via Implicit Profiles

### v2

@@ 5 Experiments
We evaluate user simulators’ authenticity and consistency at both the utterance and conversation levels while assessing diversity through the difference between our simulated and real user distributions.
@@ 5.1 Datasets
We select the popular LMSYS-Chat-1M Zheng et al. (2023) as our data source for the experiment, which contains one million human-LLM conversations. Following previous work Kong et al. (2024), we filter the samples for non-English language, toxicity, and redundancy and obtain a complete 94,874 samples (87,882/4,626/2,366 for Training/Validation/Test datasets). Then, we use a GPT-4o-based profile extractor described in Section 4.1 to preprocess them with annotating user profiles for each conversation and construct them into LMSYS-USP. Detailed preprocessing are in Appendix A.1.
@@
Table 2: Automatic evaluation results of profile quality across different datasets.
We used DPC and Val.Score to automatically evaluate the quality of extracted user profiles on the test set of LMSYS-USP, plus Persona-Chat Zhang et al. (2018) and ConvAI2  [^2] We use the human-to-bot dataset in https://huggingface.co/datasets/convai-challenge/conv_ai_2 Dinan et al. (2019) with manually annotated dialogue datasets of other existing predefined profiles. Table 2 shows that the extracted profile can achieve over 84% DPC and even the distill-llama3 is close to that of GPT-4o, demonstrating the effectiveness of our annotation. Additionally, we select 100 samples for manual evaluation and it shows over 4/5 scores for the quality of the generated profiles (See Appendix B.4 for further details).
@@ 5.2 Baseline Models
(1) User Simulator without User Profile: This includes the untrained DialogueGPT(4o), where GPT-4o relies solely on context to predict the next user utterance, and PlatoLM Kong et al. (2024), which is fine-tuned on our training dataset using LLaMA-3-8B and can be considered equivalent to our approach without incorporating profile.
(2) User Simulator Guided by the User Profile: We adopt ProfileGPT(4o) and ProfileGPT(llama), which utilizes GPT-4o and LLAMA-3-8B-Instruct AI@Meta (2024) as the role-play backbone with our constructed profiles. Additionally, we include CharacterGLM Zhou et al. (2024a), which performs role-playing with any given profile, and CharacterLLM Shao et al. (2023), which role-played on several famous celebrities. The detailed setup and prompts are provided in Appendix B.3.
@@ 5.3 Metrics
Authenticity: We use SimCSE Gao et al. (2021) to compute semantic similarity (Sem-Sim) and style embeddings Wegmann et al. (2022) to compute style similarity (Style-Sim) for evaluating $D_{\text{utt}}(u_{ij},u^{\prime}_{ij})$ and $D_{\text{dia}}(d_{i},d_{i}^{\prime})$. We also employ Author Verification Accuracy (AVA) to assess stylistic consistency by measuring whether paired sentences share authorship based on similarity thresholds Wegmann et al. (2022). For multi-turn evaluation, we compute dialogue-level distances by concatenating each user’s utterances.
Consistency: We evaluate dialogue-level consistency between dialogues and profiles using reverse metrics: r-DP.P and r-DP.R, calculated as $Fact.Con(P_{i},d_{i}^{\prime})$ and $Fact.Con(d_{i}^{\prime},P_{i})$ respectively, along with their harmonic mean r-DPC, from a profile-centric perspective. For utterance-level evaluation, we compute the mean DP.P score across individual utterances. Additionally, we incorporate Persona Coverage (P.Cover) Song et al. (2019) to assess keyword-level consistency. The Subjective Characteristic Score (SC.Score), measures the reflection of subjective traits, assessed by GPT-4o with prompt in Appendix D.
Diversity: We compute the Absolute Difference Value (ADV), which represents the Euclidean distance between the PCA-reduced embeddings of generated and target dialogues, to evaluate the discrepancy between the distribution of the reconstructed dialogues and the original dialogues.
Additionally, we assess multi-turn dialogue continuity using the early stop rate (ESR), which flags premature endings caused by repetitive responses or repeated gratitude expressions over three turns.
@@ 5.4.1 Utterance-Level Evaluation
ProfileGPT(llama) 39.82 14.88 13.47 82.19 72.29 3.92
ProfileGPT(4o) 41.66 5.74 9.87 92.73 73.34 4.71
USP w/o RLCC 54.25 46.57 43.61 71.30 71.56 3.36
@@
Table 3: Utterance-level performance comparison of different models on authenticity and consistency metrics.
@@
ProfileGPT(llama) 31 46.84 10.58 11.63 67.09 29.98 41.44 47.72 4.19
ProfileGPT(4o) 32 48.87 10.15 11.26 76.59 43.72 55.66 51.02 4.56
USP w/o RLCC 12 66.17 40.01 35.68 53.17 71.88 61.13 42.63 3.24
@@
Table 4: Conversation-level performance comparison of different models on authenticity and consistency metrics.
In the utterance-level evaluation, we evaluate the quality of a single-turn response generated by the testing models in a given context.
As shown in Table  3, USP outperforms all baselines in terms of authenticity, with 53.38 and 46.60, as measured by both semantic (Sem-Sim) and stylistic (Style-Sim) similarity metrics. This highlights the effectiveness of our implicit profile-assisted approach for user-LLM dialogue reconstruction, particularly when compared to context-only models like PlatoLM. Although dedicated role-playing models (ProfileGPT variants) achieve higher consistency scores (r-DP.P), this can be attributed to their direct profile text copying. USP maintains comparable overall performance while striking a better balance between authenticity and consistency.
5.4.2 Conversation-Level Evaluation
In the conversation-level evaluation, we assess the performance of testing models to chat with GPT-4o in multi-turns, providing the profile or the first turn of the reference dialogue according to their needs.
As shown in Table 4, USP outperforms baseline models in authenticity, consistency, and continuity. With the lowest ESR(10), it demonstrates superior dialogue continuity. Notably, USP’s advantage in authenticity is more pronounced in conversation-level scenarios than sentence-level predictions. In terms of consistency, USP demonstrates exceptional performance in r-DP.R metrics and achieves significantly higher r-DPC scores for overall profile dialogue consistency. This superior performance, particularly when compared to role-playing models such as ProfileGPT(4o) and ProfileGPT(llama) which show high P.Cover, suggests that RLCC effectively captures abstract profile characteristics beyond merely matching surface-level keywords. See Appendix C.2 for further case study analysis.
@@ 5.4.3 Human Evaluation
We also randomly selected 100 samples with 8 evaluators to conduct conversation-level human evaluations considering authenticity and consistency. Authenticity was assessed through Style, Semantics, and Quality, while consistency covered Accuracy, Completeness, and Quality. Full evaluation details are provided in Appendix B.4.
Table 5 demonstrates USP’s superior performance in both authenticity and consistency metrics. Our USP significantly outperforms ProfileGPT (4o) in terms of authenticity (74 vs. 13) and consistency (61 vs. 35) in manual evaluation. USP is superior to PlatoLM trained on the same dataset in terms of authenticity, which demonstrates the usefulness of implicit profile modeling. Thanks to the RLCC module, our USP model has significantly improved consistency (43 vs. 30) by aligning user profiles.
@@
USP w/o RLCC 37/32/31 43/27/30
Table 5: Human evaluation results comparing baselines with USP on authenticity and consistency.
@@ 5.4.4 Diversity Sampling Evaluation
Table 6: Ablation study of hyperparameters in RLCC.
Figure 3 shows the absolute difference value between target dialogue and generated dialogue by various models across different percentiles. From the results, we observe that USP and USP w/o RLCC consistently achieve the smallest ADV across all percentiles, indicating the dialogues they generated that closely match the target conversations. For example, marked by the red cross, PlatoLM has 60% of sample ADV below 15%, while USP has only 5% or less ADV. It demonstrates its stronger capability in preserving the semantic characteristics of the original dialogues. The uniformly lower curves of USP and USP w/o RLCC compared to other baselines (PlatoLM, ProfileGPT(4o), and DialogueGPT(4o)) suggest that our approach generates dialogues that are more faithful to the target conversations across typical and extreme cases.
We also analyze and demonstrate that our user simulator can sample different representatives (majority and minority) of users compared to random sampling in Appendix C.1 and apply it to downstream applications in Appendix C.3.
Figure 3: Cumulative distribution of ADV performance comparison across different models.
6 Analysis
6.1 Ablation Study
To evaluate the relative importance of RLCC’s two rewards, we tested different values of $\lambda$ in Equation 6, denoted as USP($\lambda$: $1-\lambda$).
As shown in Table 6, $\lambda$ = 0.8 provides the optimal balance between maintaining model capabilities and enhancing dialogue consistency. Higher values of $\lambda$ (0.9) compromise speaking style authenticity without improving r-DPC, leading to superficial profile matching, as shown by increased P.Cover scores. Conversely, $\lambda$ = 0.5 achieves authentic style features but lacks sufficient consistency emphasis, resulting in stagnant performance across capabilities.
@@ 6.2 Applications: Dynamic Multi-turn Evaluation For LLMs
One application of our simulator is to fill the gap in the current dynamic multi-turn evaluation of large models. The user simulator can simulate different user groups to dynamically interact with the tested model in multiple rounds and reveal their specific defects, as shown in Table 7.
@@
Table 7: Response model performance comparison over different target groups.
We simulated 300 diverse user profiles by USP, comprising 100 highest-probability profiles(the majority), 100 lowest-probability profiles(the minority) based on estimated density, and 100 random synthetic profiles(the virtuality), using the sampler mentioned in Section 4.4. Then, we leverage USP using these profiles to chat with LLM in multi-turns and evaluate them, followed by MT-Bench (Zheng et al., 2024). The experimental results indicate that our user simulator is effective and consistent with the average rankings on the latest Livebench White et al. (2024) and Chatbot-Arena Chiang et al. (2024). In addition, it can also be seen that GPT-4o is better at catering to minority groups than other models, indicating its superior robustness. A more detailed analysis can be found in the appendix C.3.
@@ 7 Conclusion
In this work, we propose a novel user simulator with implicit profiles that excels in authenticity, consistency, and diversity. Based on this, we introduce the USP framework, which integrates extracted user profiles into the user simulator by conditional fine-tuning and reinforcement learning with cycle consistency. Our experimental results, validated by both automatic metrics and human evaluations, show that USP significantly outperforms role-playing simulators (e.g., GPT-4o) and direct simulation approaches (e.g., PlatoLM) in authenticity while achieving comparable consistency at both the sentence and conversation levels. Furthermore, through a dynamic evaluation across various LLMs chatting with diverse demographic groups, we demonstrate USP’s effectiveness in real-world applications.
@@
We acknowledge the following limitations: 1) Applicability Across Different Scenarios: We conduct the experiments on a single dataset, and there has been limited validation across multiple datasets to assess the generalizability of the results. 2) Cultural and Linguistic Scope: We focus on English dialogues in this paper, which may limit the applicability of USP to other linguistic and cultural contexts.
Ethics Statement
Although LMSYS-1M has undergone extensive data cleaning and ethical checks, the dataset may still contain sensitive or harmful content, reflecting violent, explicit, or discriminatory traits in certain dialogue. This could result in USP generating unsafe dialogues. We strongly advise against including sensitive terms in profiles when using USP, as this may lead to extreme behavior in both the USP and the response model.
@@ A.2 Profile Dataset
As discussed in Section 4.1, we categorize all attributes into three types: scene-consistent attributes, scene-related attributes, and deep intrinsic characteristics. For scene-consistent attributes, we use the prompt shown in Figure 13, with each metric definition following the guidelines outlined in Cheng et al. (2024b). For scene-related attributes, we use the prompt in Figure 14, and for deep intrinsic characteristics, we refer to Figure 12. The definition of the Big Five Traits scores follows Gosling et al. (2003).
Next, we concatenate the attributes, remove invalid values (e.g., null or meaningless values), and shuffle the order to eliminate any positional bias in the generated profiles. The attributes, which encompass the three aforementioned aspects, are then rephrased using GPT-4o, with the prompt shown in Figure 15. As a result, we obtain automatically labeled profiles for each data entry. The length statistics are shown in Table 8.
Dataset Train Val Test Profile
LMSYS-USP 1,149 1,295 1,438 231
Table 8: Average token length of LMSYS-USP dataset.
Furthermore, we calculated the frequency of occurrence for each attribute value (i.e., the average number of different attribute values per sample) to assess the prevalence of each attribute. The statistics for the objective facts can be found in Figure 4, while for the subjective characteristics, we focused on whether the Big Five Traits were significantly exhibited. Specifically, we only consider traits with high or low scores, while moderate scores are viewed as the average representation of human behavior Moon et al. (2024) and are not included in the subsequent profiles.
@@
Figure 4: Frequency of occurrence of values across different attributes of objective facts in the attribute extraction process.

### v3

@@ 5 Experiments
We evaluate user simulators on authenticity, consistency, and multi-turn continuity at both utterance and conversation levels, while measuring diversity by comparing the dialogue distributions of simulated and real users.
@@ 5.1 Datasets
We select the popular LMSYS-Chat-1M Zheng et al. (2023), which contains one million human-LLM conversations. Following prior work Kong et al. (2024), we filter out non-English, toxic, and redundant samples, resulting in 94,874 conversations (87,882 for training, 4,626 for validation, and 2,366 for testing). Each conversation is then annotated with user profiles using the GPT-4o-based extractor described in Section 4.1, forming the LMSYS-USP dataset. Detailed preprocessing steps are provided in Appendix A.1.
@@
Table 2: Automated evaluation of profile quality across datasets. Avg DP.P # Fact denotes the average number of atomic facts per user profile, while Avg DP.R # Fact represents the average number of user utterances per dialogue. Note that human-annotated profiles in PersonaChat and ConvAI2 contain no subjective characteristics.
We use DPC and Val.Score to automatically evaluate the quality of extracted user profiles on 100 randomly selected samples from the LMSYS-USP test set, as well as on 100 samples each from Persona-Chat Zhang et al. (2018) and ConvAI2111We use the human-to-bot dataset from https://huggingface.co/datasets/convai-challenge/conv_ai_2 Dinan et al. (2019), which include manually annotated predefined profiles. As shown in Table 2, the extracted profiles achieve over 84% DPC, with even distill-llama3 results comparable to GPT-4o, demonstrating the effectiveness of our annotation method. Manual evaluation further confirms profile quality, with average scores exceeding 4 out of 5 (see Appendix B.3 for details).
5.2 Configurations
We train USP based on LLaMA-3-8B-Base AI@Meta (2024) model. The conditional SFT is conducted on the training dataset using 4 A100 40GB GPUs, with full fine-tuning over 3 epochs at a learning rate of 5e-5 and max length set to 4096, taking approximately two days. Our diverse profile sampler then randomly selects 1,000 samples from the training set for virtual user sampling, combining objective facts and subjective characteristics to generate about 1 million profiles. From these, we select the 5,000 profiles least similar to the training dataset for the RLCC phase. RLCC training uses two H20 96GB GPUs for 5 days, utilizing a KL coefficient of 0.01, a learning rate of 5e-7, and training for 1 epoch.
@@ 5.3 Baseline Models
(1) User Simulator without User Profile: This includes untrained GPT-4o (User w/o Profile), which use GPT-4o to predict user utterances based solely on context, and PlatoLM Kong et al. (2024), a baseline fully fine-tuned on our dataset using LLaMA-3-8B-Base, representing a profile-agnostic approach.
(2) User Simulator with User Profile: We employ GPT-4o (User w/ Profile) and LLaMA3 (User w/ Profile) leveraging GPT-4o and LLaMA-3-8B-Instruct AI@Meta (2024) as profile-conditioned role-playing agents, alongside CharacterGLM Zhou et al. (2024a), a flexible profile-based baseline, and CharacterLLM Shao et al. (2023), designed to emulate public figures.
@@ 5.4 Metrics
Authenticity: We evaluate semantic and stylistic similarity using SimCSE Gao et al. (2021) and style embeddings Wegmann et al. (2022), respectively, to compute $D_{\text{utt}}(u_{i,j},u^{\prime}_{i,j})$ and $D_{\text{dia}}(d_{i},d_{i}^{\prime})$. To assess stylistic consistency, we report Author Verification Accuracy (AVA) Wegmann et al. (2022), which measures whether sentence pairs are attributed to the same author based on similarity thresholds. Dialogue-level distances are computed by concatenating all user utterances.
[…]
Diversity: We measure the Absolute Difference Value (ADV), defined as the Euclidean distance between PCA-reduced embeddings of generated and target dialogues, to quantify the distributional discrepancy between simulated and target dialogues.
Continuity: Multi-turn dialogue continuity ability is evaluated via the early stop rate (ESR), which detects premature endings triggered by repetitive responses or repeated expressions of gratitude across three turns.
@@ 5.5.1 Utterance-Level Evaluation
LLaMA3 (User w/ Profile) 39.82 14.88 13.47 82.19 72.29 3.92
GPT-4o (User w/ Profile) 41.66 5.74 9.87 92.73 73.34 4.71
USP (w/o RLCC) 54.25 46.57 43.61 71.30 71.56 3.36
@@
Table 3: Utterance-level performance comparison of different user simulator.
In utterance-level evaluation, we assess the quality of single-turn responses generated by different user simulators given the golden context.
As shown in Table 3, USP outperforms all baselines in authenticity, as measured by both semantic similarity (Sem-Sim: 53.38) and stylistic similarity (Style-Sim: 46.60). This shows the effectiveness of our implicit profile-based approach for user-LLM dialogue reconstruction, especially compared to non-profile baselines like PlatoLM. While dedicated role-playing models (e.g., GPT-4o (User w/ Profile)) achieve higher consistency scores (r-DP.P) due to direct profile keyword copying with high P.Cover (73.34), our USP strikes a better balance between authenticity and consistency, as shown by the intuitive examples in Section 6.3.
5.5.2 Conversation-Level Evaluation
@@
LLaMA3 (User w/ Profile) 31 46.84 10.58 11.63 67.09 29.98 41.44 47.72 4.19
GPT-4o (User w/ Profile) 32 48.87 10.15 11.26 76.59 43.72 55.66 51.02 4.56
USP (w/o RLCC) 12 66.17 40.01 35.68 53.17 71.88 61.13 42.63 3.24
@@
Table 4: Conversation-level performance comparison of different user simulators.
In the conversation-level evaluation, we assess the quality of multi-turn dialogues generated by different user simulators interacting with GPT-4o, each provided with either a given profile or the first turn of a reference dialogue.
As shown in Table 4, USP outperforms baseline models in authenticity, consistency, and continuity. With the lowest ESR (10), USP ensures superior dialogue continuity, avoiding issues like repetitive generation and reciprocal appreciation loops seen in baselines. Its advantage in authenticity is especially evident in multi-turn scenarios, compared to sentence-level evaluations. In terms of consistency, USP excels with a high r-DP.R (74.38) and significantly better r-DPC (64.05), demonstrating strong conditional generation consistency. Unlike role-playing models such as GPT-4o (User w/ Profile) and LLaMA3 (User w/ Profile), which show high P.Cover and r-DP.P but lower overall profile dialogue consistency, USP demonstrates a deeper and more comprehensive understanding of user behavior, moving beyond surface-level keyword matching to deliver a more vivid user simulation.
@@ 5.5.3 Human Evaluation
We randomly selected 100 samples from the test set and engaged 8 evaluators to assess conversations on authenticity and consistency. Authenticity was evaluated based on Style, Semantics, and Quality, while consistency focused on Accuracy, Completeness, and Quality. Detailed criteria are in Appendix B.3.
Table 5 shows USP’s clear superiority in both authenticity and consistency. USP outperforms GPT-4o (User w/ Profile) in authenticity (74 vs. 13) and consistency (61 vs. 35). It also surpasses PlatoLM trained on the same data in authenticity, demonstrating the advantage of implicit profile modeling. The larger gap in consistency (43 vs. 30) compared to authenticity (37 vs. 31) between USP and USP (w/o RLCC) highlights RLCC’s key role in aligning profiles with dialogues.
@@
USP (w/o RLCC) 37/32/31 43/27/30
Table 5: Human evaluation of USP win rates over baselines in terms of authenticity and consistency. $\kappa$ denotes the within-group kappa coefficient. Note that PlatoLM, as a non-profile baseline, contain no consistency results.
@@ 5.5.4 Diversity Sampling Evaluation
Figure 3: Cumulative distribution of ADV performance comparison across different simulators, with the red cross indicating that USP (w/o RLCC) has 60% of its samples with ADV below 5%.
Figure 3 compares ADV between target and generated dialogues across simulators and percentiles. The USP series consistently demonstrate lower ADV, even in extreme cases, with 60% of samples achieving ADV below 5% (marked by a red cross), compared to baselines(e.g. PlatoLM, GPT-4o (User w/ Profile)) at 10% or higher. This reflects the USP series’ superior ability to generate dialogues that closely align with target conversations, effectively preserving the diversity distribution of user characteristics.
We also show that our sampling strategy outperforms random sampling by effectively capturing diverse representatives (majority and minority) in Appendix C.1.
6 Analysis
6.1 Ablation Study
@@
Table 6: Ablation study of our USP framework.
We evaluate the effectiveness of the polishing step in our two-stage profile construction pipeline, which converts attributes into natural language descriptions, by comparing it to a baseline (USAP w/o RLCC) that uses only attributes without polishing. As shown in the first two rows of Table 6, the polishing step enhances generalization, improving performance across most metrics of Continuity, Authenticity, and Consistency. In contrast, relying solely on attributes leads USAP (w/o RLCC) to excessively replicate profile descriptions, resulting in a high P.Cover score (50.35) due to attributes appearing directly in the dialogue.
We also assess the relative importance of RLCC’s two rewards by testing different $\lambda$ values in Equation 5, denoted as USP($\lambda$: $1-\lambda$). Table 6 shows that $\lambda=0.8$ optimally balances model capabilities and dialogue consistency. Higher $\lambda$ (0.9) sacrifices speaking style authenticity without improving r-DPC, increasing P.Cover and indicating superficial profile matching. Conversely, $\lambda=0.5$ preserves authentic style but lacks sufficient consistency, resulting in stagnant performance.
@@ 6.2 Applications: Dynamic Multi-turn Evaluation For LLMs
One application of our simulator is addressing the gap in dynamic multi-turn evaluation of LLMs. While current automatic evaluations rely on static preset questions Zheng et al. (2024); Bai et al. (2024), our user simulator can dynamically interact with the LLM over multiple rounds, adjusting based on response quality and given user traits.
We generated 300 diverse user profiles using our sampler: 100 highest-probability (majority), 100 lowest-probability (minority), and 100 random synthetic (virtual) profiles. Using these profiles, USP engages in multi-turn dialogues with the LLM. Evaluation results, based on MT-Bench (Zheng et al., 2024) and presented in Table 7, show that dynamic multi-turn evaluation aligns closely with average rankings on LiveBench White et al. (2024) and Chatbot-Arena Chiang et al. (2024), while revealing fine-grained weaknesses of different LLMs when interacting with specific user groups. Detailed analysis is in Appendix C.3.
@@
Table 7: LLM performance across user groups.
6.3 Case Study
Table 8 shows that the role-playing baseline GPT-4o (w/ Profile) often copies abstract profile traits verbatim to assert user identity. In contrast, our USP conveys these traits more naturally by transforming abstract concepts into concrete and coherent expressions. For instance, when the profile states “being a father of two,” GPT-4o (w/ Profile) repeats it directly, while USP implicitly reflects this by mentioning a “son” and a “daughter” in later turns. Similarly, rather than restating “likes Italian food,” USP refers to a specific dish like pasta. These examples illustrate that USP better mimics human behavior by expressing abstract, high-level traits both directly and subtly in a natural manner, likely contributing to its stronger human preference (see Table 5). Additional case studies are provided in Appendix C.2.
(Profile) As a father of two, your strong sense of family responsibility is evident… Your interest in vegan Italian-inspired snack dishes
(1st turn) GPT-4o (w/ Profile): As a father of two who has a keen interest in …
(2nd turn) GPT-4o (w/ Profile): Lately, I’ve been exploring new vegan Italian-inspired snack dishes
(4th turn) USP:My son wants pasta but he doesn’t like tomato sauce…
(5th turn)my daughter says she does not feel well… She prefers vegan food in Italian style
Table 8: Case study comparing outputs from USP and the GPT-4o (w/ Profile) baseline for a sample profile. Bold highlights keywords explicitly copied by GPT-4o, while bold italic marks USP’s implicit or fuzzy matches.
@@ 7 Conclusion
In this work, we introduce the USP framework, which integrates extracted user profiles into the user simulator by conditional SFT and RLCC. Our experimental results, validated by both automatic metrics and human evaluations, show that USP significantly outperforms role-playing simulators (e.g., GPT-4o (User w/o Profile)) and direct simulation approaches (e.g., PlatoLM) in authenticity and diversity while achieving comparable consistency at both the sentence and conversation levels. Additionally, dynamic evaluations with various LLMs across diverse demographic groups highlight USP’s effectiveness in real-world scenarios. Nonetheless, a gap remains compared to real human behavior, and our future work will explore finer-grained control and multimodal simulation.
@@
We acknowledge the following limitations: 1) Scenario Applicability: Experiments were conducted on a single dataset, with minimal validation across others to confirm broader applicability. 2) Linguistic and Cultural Scope: Our focus on English dialogues may limit the applicability of USP to other languages and cultural contexts.
Acknowledgments
This research is supported by the project of Shenzhen Science and Technology Research Fund (Fundamental Research Key Project, Grant No. JCYJ20220818103001002), Shenzhen Science and Technology Program (Shenzhen Key Laboratory, Grant No. ZDSYS20230626091302006), Shenzhen Stability Science Program 2023, Shenzhen Key Lab of Multi-Modal Cognitive Computing, SRIBD Innovation Fund (Grant No. K00120240006), and Program for Guangdong Introducing Innovative and Entrepreneurial Teams, Grant No. 2023ZT10X044.
@@ A.2 Profile Dataset
As detailed in Section 4.1, we classify attributes into three types: scene-consistent, scene-related, and deep intrinsic characteristics. For each, we use specific prompts (Figures 11, 12, and 13), with metric definitions based on Cheng et al. (2024b) and Big Five traits per Gosling et al. (2003).
We then concatenate these attributes, remove invalid entries, and shuffle their order to prevent positional bias. The combined attributes are rephrased using GPT-4o with the prompt in Figure 14, producing automatically labeled profiles. The LMSYS-USP dataset averages 1,149 tokens in training, 1,295 in validation, 1,438 in testing, and 231 tokens per profile.
We also measured the frequency of each attribute value, defined as the average number of distinct values per sample, to assess attribute prevalence. Statistics for objective facts are shown in Figure 4. For subjective traits, we focused on the Big Five traits only when scores were significantly high or low, excluding moderate scores as they reflect average human behavior Moon et al. (2024) and were omitted from the subsequent polishing step.
@@
Figure 4: Frequency of values for each attribute of objective facts in the attribute extraction process.